9. The number of images formed by two plane mirrors placed at angle 90 degree is
a.four
b.three
c.two
d.one

10.The number of images formed by two plane mirrors placed at angle 180 degree is
a.two
b.one
c.three
d.four
1
Expert's answer
2019-03-19T12:20:20-0400

The number of images formed by two plane mirrors placed at angle A degree is

"n=\frac{360}{A}-1"

If A=90 degree

"n=\frac{360}{90}-1=3"

If A=180 degree

"n=\frac{360}{180}-1=1"